# Covariant tomography of fields

### v1

@@ Abstract
This paper addresses the Inverse Boundary Value Problem (IBVP) for classical fields, specifically focusing on the recovery of parallelly transformed fields within a region based on known boundary data. We introduce a local solution framework, termed ”covariant tomography,” that uses geometric decomposition to reconstruct interior fields and currents within star-shaped open subsets. The core of our approach involves decomposing differential forms into exact and antiexact components, enabling the formulation of the parallel transport equation via a homotopy operator. We examine three primary extension techniques - radial, heat equation, and harmonic - to map boundary values into the interior, noting that the choice of extension directly influences the regularity of the resulting currents. The proposed methodology provides a systematic way to identify the realizability of boundary values and offers solutions for both current and gauge field tomography. Finally, we demonstrate the utility of this framework through illustrative examples in low-dimensional spaces and electromagnetic potential reconstruction in $\mathbb{R}^{3}$.
@@ 1 Introduction
The specific problems for ODEs, finite difference, and integral equations [5], PDEs [33], [13], or, in particular, exterior differential equations [32] can be defined.
@@ 3.1 Extension problem
    Constant extension along the rays.
@@
For extension by the constant value at the boundary along the ray we have to use the homotopy $x=F(x_{b},t_{x})=x_{0}+t_{x}(x_{b}-x_{0})$ that connects $x_{b}\in B$ with the center $x_{0}$ and such that $x$ is on the ray $x_{0},x_{b}]$ for the homotopy parameter $t_{x}\in[0,1]$. We assume that $x$ is on the ray.
@@
The exterior derivative in $d^{\nabla}$ in this case must be understood as a distributional derivative. In the case of discontinuity of $\Phi$, the current $J_{ext}$ is a singular distribution that compensates it. It also measures the non-covariant constancy of $\Phi$.
@@
The diffusion problem generally smooths the initial conditions instantaneously [30]. The current (36) is now smooth, and its nonzero values indicate compensation currents of extension.
The last case is formally the previous case with $t\rightarrow\infty$. More precisely, as before, we define a metric structure that results in the Laplace-Beltrami operator. Then the extension problem is equivalent to solving
@@
Knowing the solution $\Phi$, the extension operator $\tilde{\Phi}(\alpha)=\Phi$. As harmonic solutions are smooth, like for the heat equation, the extension current (36) will be regular.
@@ 3.2 Projection problem
Case - $A$ fixed: As for first case assume that the connection form $A$ is fixed. Then $d^{\nabla}$ is also fixed. We compute the current $J_{ext}=d^{\nabla}\Phi$. We consider three cases.
@@
Case - $J$ fixed: When $J$ is fixed we have to solve the problem for $A$ that yields algebraic equations of the form
@@ 3.4 Tomography for geometric-based equations
These two operators can be composed in multiple ways, yielding so-called geometric-based differential equations. As discussed in [22], the equation containing a composed operator can be decomposed into covariant and contravariant parallel transport equations exactly in the same way as higher-order ODEs can be rewritten as a system of first-order ODEs. As an example, consider
@@
can be decomposed as
$\reflectbox{D}\phi_{1}=\phi_{2},$ (52)
$D\phi_{2}=J.$ (53)
To show the IVBP for (51), we adjoint the condition $j^{*}\phi=\alpha$. Then, using the decomposition, we have the system
@@ 4 Conclusions
In this work, we have established a robust framework for solving Inverse Boundary Value Problems related to the parallel transport of differential forms. By leveraging geometric decomposition, we successfully bridged the gap between boundary restrictions and interior field configurations, providing clear criteria for recovering currents and gauge potentials.
A significant advancement of this research is the extension of the tomography framework to geometric-based differential equations, which incorporate both the exterior derivative and the codifferential to address complex, higher-order systems like the Maxwell equations. While the method offers an efficient local solution through formal series expansion, we have identified critical constraints regarding the radius of convergence and the non-uniqueness inherent in gauge modes. These results demonstrate that the application of the homotopy operator and the Poincaré lemma facilitates a powerful, algorithmically straightforward approach to IBVP. This research opens a new way for applying exterior calculus to tomographic problems in physics and engineering, where the underlying geometry plays an essential role in field reconstruction.
@@ Acknowledgments
The author acknowledges the support of the COST CA24122 action.

### v2

# arXiv:2601.13261v2[math-ph] 27 Feb 2026

@@ Abstract
This paper develops ’covariant tomography’, a local framework for solving Inverse Boundary Value Problems (IBVP) for parallel transport equation on star-shaped domains. By integrating geometric decomposition with specific interior extensions - radial, heat equation, or harmonic - the method reconstructs currents and gauge potentials from boundary data. The choice of extension directly dictates the regularity of the recovered interior fields. A primary contribution is the ’tower’ algorithm, which reduces higher-order systems, such as Maxwell equations, to a sequence of coupled first-order equations. We establish a formal solvability criterion (Theorem 6), proving that higher-order IBVPs are solvable if and only if this tower is sequentially solvable. The framework is validated through low-dimensional examples and electromagnetic potential reconstruction in $\mathbb{R}^{3}$.
@@ 1 Introduction
Historically, the problem was formulated in [14] for the inverse Sturm-Liouville problem and pursued in [32] for the multidimensional inverse problem for the Schrödinger equation. The specific problems for ODEs, finite difference, and integral equations [5], PDEs [37], [16], or, in particular, exterior differential equations [36] can be defined.
@@
  • Electromagnetic problems can be treated by Inverse Scattering Problem [8, 6].
@@ 3.1 Extension problem
    Constant extension along the rays - radial extension.
@@
For extension by the constant value at the boundary along the ray - radial extension - we have to use the homotopy $x=F(x_{b},t_{x})=x_{0}+t_{x}(x_{b}-x_{0})$ that connects $x_{b}\in B$ with the center $x_{0}$ and such that $x$ is on the ray $x_{0},x_{b}]$ for the homotopy parameter $t_{x}\in[0,1]$. We assume that $x$ is on the ray.
@@
The exterior derivative in $d^{\nabla}$ in this case must be understood as a distributional derivative. In the case of discontinuity of $\Phi$, the current $J_{ext}$ is a singular distribution that compensates it. It also measures the non-covariant constancy of $\Phi$. We obtain
Proposition 2.
The radial extension maps $C^{k}(B)$ boundary data onto $C^{0}(U\setminus\{x_{0}\})$ interior solution.
@@
The diffusion problem generally smooths the initial conditions instantaneously [34]. The current (36) is now smooth, and its nonzero values indicate compensation currents of extension. Considering the heat equation property we obtain
Proposition 3.
The heat extension maps $L^{2}(B)$ boundary value onto $C^{\infty}(U\times(0,T])$ interior solution.
The last case - harmonic extension - is formally the previous case with $t\rightarrow\infty$. More precisely, as before, we define a metric structure that results in the Laplace-Beltrami operator. Then the extension problem is equivalent to solving
@@
Knowing the solution $\Phi$, the extension operator $\tilde{\Phi}(\alpha)=\Phi$. As harmonic solutions are smooth when starting from smooth boundary data, the extension current (36) will be regular. We have
Proposition 4.
The harmonic extension maps $C^{\infty}(B)$ boundary value onto $C^{\infty}(U)$ interior solution.
@@ 3.2 Projection problem
Case - $A$ fixed: We have the following:
Theorem 4.
Let $U$ be a star-shaped subset and $A$ a fixed connection one-form. Given boundary data $\alpha$, a unique interior current $J$ can be recovered (up to gauge mode) trough the mapping $J=d^{\nabla}\Phi$, where $\Phi$ is the harmonic extension of $\alpha$.
Proof.
The connection form $A$ is fixed, therefore $d^{\nabla}$ is also fixed. We compute the current $J_{ext}=d^{\nabla}\Phi$. We consider three cases.
@@
Case - $J$ fixed: We have:
Theorem 5.
Given an interior current $J$ and an extended field $\Phi$ form the boundary field $\alpha$, the connection $A$ is (non-uniquely) determined by the algebraic relation $A\wedge\Phi=J-d\Phi$.
Proof.
When $J$ is fixed we have to solve the problem for $A$ that yields algebraic equations of the form
@@ 3.4 Tomography for geometric-based equations
These two operators can be composed in multiple ways, yielding so-called geometric-based differential equations. As discussed in [25], the equation containing a composed operator can be decomposed into covariant and contravariant parallel transport equations exactly in the same way as higher-order ODEs can be rewritten as a system of first-order ODEs. This is provided by the following:
Algorithm 1 (Tower Method).
In order to solve the geometric-based equations of the form
$\left\{\begin{array}[]{c}D\phi=J,\\ j^{*}\phi=\alpha\end{array}\right$. (51)
with a fixed boundary data $j^{*}\phi=\alpha$, where $D$ is the $k$ composition of $D_{A}$ and $\reflectbox{D}_{X}$ operators
  1. 1.
    Decompose the equation into a system of coupled first-order equations
    $\left\{\begin{array}[]{c}D_{k}\phi_{k}=J\\ D_{k-1}\phi_{k-1}=\phi_{k}\\ \ldots\\ D_{2}\phi_{2}=\phi_{3}\\ D_{1}\phi_{1}=\phi_{2}\\ j^{*}\phi_{k}=\alpha,\end{array}\right$. (52)
    where $D_{i}$ is either $D_{A}$ or $\reflectbox{D}_{X}$ operator.
  2. 2.
    Solve the IBVP $D_{k}\phi_{k}=J$, $j^{*}\phi_{k}=\alpha$ as described in the previous section.
  3. 3.
    Propagate the solution $\phi_{k}$ for other $\phi_{i}$ solving recursively remaining parallel/covariant transport equations.
We therefore have
Theorem 6.
For the IBVP problem (51) to be solved if and only if associated tower of first-order equations (52) is sequentially solvable.
We illustrate the algorithm 1 by simple example
Example 1.
Consider
@@
$j^{*}\phi=\alpha$ (54)
which can be decomposed as
@@ 3.5 Examples
In the next examples we consider IBVP for Maxwell equations [31, 42].
@@ 4 Conclusions
This research establishes a robust framework for solving Inverse Boundary Value Problems (IBVPs) through a methodology termed ’covariant tomography’. By integrating geometric decomposition with the linear homotopy operator, we have successfully bridged the gap between boundary restrictions and interior field configurations. This approach provides a systematic way to recover interior currents and gauge potentials within star-shaped domains.
A significant novelty of this work is the introduction of Algorithm 1, which decomposes higher-order, geometric-based differential equations into a ’tower’ of coupled first-order systems. This allows for the tomographic reconstruction of fields, such as solutions of the Maxwell equations, through sequential parallel transport equations. This framework is formally supported by Theorem 6, which establishes a solvability criterion: a higher-order IBVP is solvable if and only if its associated first-order tower is sequentially solvable.
While the method offers an algorithmically straightforward local solution via formal series expansion, it is subject to specific constraints:
  • Convergence: The solution is limited by a finite radius of convergence dependent on the connection norm.
  • Non-uniqueness: The presence of gauge modes inherent in the geometric decomposition can lead to non-unique interior configurations.
  • Regularity: The smoothness of recovered currents is directly influenced by the choice of extension technique-radial, heat equation, or harmonic.
By grounding these results in the historical progress in multidimensional inverse problems - referencing the foundational work of Gelfand [14] and Novikov [32] - this research opens new path for applying exterior calculus to field reconstruction in physics and engineering.
@@ Acknowledgments
The author acknowledges the support of the COST CA24122 and CA22123 actions.